Find the minimum value of n in the Balmer series for which the predicted wavelength is in the ultraviolet region of the spectrum.
Answer #1

λ=B*(n^2)/(n^2 - 2^2)
λ=(B*n^2) / (n^2 - 4)
λ*(n^2 -4)=B*n^2
n^2*(λ-B)=4λ
n^2=4λ/(λ-B)
n=sqrt (4λ/(λ-B))

The minimum n is for the longest ultraviolet wavelength which is 4*10^-7m=400nm
Therefore:

n=sqrt(4*4*(10^-7)/(4*(10^-7)-3.6456×(…
n=6.71913...

Therefore the minimum value of n is 7. :)